# Fully stochastic trust-region methods with Barzilai-Borwein steplengths

### v1

@@ Abstract
We investigate on stochastic gradient methods and stochastic counterparts of the Barzilai-Borwein steplengths and their application to finite-sum minimization problems. Our proposal is based on the Trust-Region-ish (TRish) framework introduced in [F. E. Curtis, K. Scheinberg, R. Shi, A stochastic trust region algorithm based on careful step normalization, Informs Journal on Optimization, 1, 2019]. The new framework, named TRishBB, aims at enhancing the performance of TRish and at reducing the computational cost of the second-order TRish variant. We propose three different methods belonging to the TRishBB framework and present the convergence analysis for possibly nonconvex objective functions, considering biased and unbiased gradient approximations. Our analysis requires neither diminishing step-sizes nor full gradient evaluation. The numerical experiments in machine learning applications demonstrate the effectiveness of applying the Barzilai-Borwein steplength with stochastic gradients and show improved testing accuracy compared to the TRish method.
@@ 1 Introduction
Computing the gradient $\nabla f(x_{k})$ to perform the update (2) is practically inconvenient in ML and DL applications where $f$ represents a loss function on a given dataset and $N$ is large. The Stochastic Gradient (SG) method [4, 5] and the mini-batch SG method [6] are iterative procedures where the full gradient $\nabla f(x_{k})$ is replaced by either a single gradient $\nabla f_{i}(x_{k})$, $i\in\mathcal{N}$ chosen randomly, or by a mini-batch of samples of the form $\nabla f_{\mathcal{N}_{k}}(x_{k})=\frac{1}{|\mathcal{N}_{k}|}\sum_{i\in\mathcal{N}_{k}}\nabla f_{i}(x_{k})$ where the set $\mathcal{N}_{k}\subseteq\mathcal{N}$ is chosen randomly and $|\mathcal{N}_{k}|$ denotes its cardinality, i.e., the mini-batch size. A limitation of such methods is the need to choose problem-dependent steplengths to guarantee convergence. An attempt to reduce such dependence was made in [7] where the steplenghts depend on a Trust-Region-ish (TRish) rule and normalized steps are selected whenever the norm of the stochastic gradient is within a prefixed interval whose endpoints need to be tuned. TRish method was designed for both first-order models [7, 8] and quadratic models [9] and, contrary to a standard trust-region approach, does not test for actual versus predicted reduction and thus does not call for $f$-evaluations. Results from [7, 8] show that TRish can outperform a standard stochastic gradient approach and it was very recently extended to equality-constrained problems [10].
@@
This paper is organized as follows. The remainder of this section is devoted to a review of the relevant literature providing background, motivation, and contribution to our work. We outline the proposed stochastic algorithms in Section 2 and present the theoretical analysis in Section 3. In Section 4 we present he implementation details, experimental setup, and an extensive evaluation of the method’s performance. Finally, in Section 5 we provide an overview of the obtained results and perspectives.
@@ 1.1 Related works and motivations
In [11], BB steplengths were incorporated into the Stochastic Variance Reduction Gradient (SVRG) method [23] and into the SG method. The proposed methods are inner-outer loop-based frameworks. Since SVRG provides a full gradient at each outer loop, the BB steplength was computed at outer loops using the standard formula for the BB steplength; convergence of the resulting procedure was proved for strongly convex objective functions. On the other hand, single stochastic gradients $\nabla f_{i}$ were employed at each iteration in the SG method and, in the inner loop, a moving averaging rule over $N$ stochastic gradients at different points was computed; then, such average was used at each outer loop to evaluate a stochastic BB parameter. A theoretical support for this method was not provided. In [12], deterministic BB steplengths from [24], based on exact gradient and function evaluations, were used in the SVRG method; convergence was analyzed for strongly convex objective functions. In [13], stochastic BB steplengths were formed by a moving averaging rule on mini-batch gradients evaluated at different points and selected if they lied within an interval with endpoints tending to zero. Using diminishing stepsizes provides convergence due to well-known results for SG method but, in principle, BB steplengths may be discarded eventually since they are not supposed to tend to zero. A further method in which stochastic BB steplengths were formed and diminishing stepsizes were imposed is given in [14]; the distinguish feature is that the mini-batch of sample is taken constant for a prefixed number of iterations with the aim to explore the underlying approximate Hessian spectrum. Finally, the paper [15] investigates on the use of BB steplenghts combined with projected stochastic gradients for constrained optimization problems and with the objective function in the form of mathematical expectation.
@@ 2 New Stochastic Trust-Region-ish Algorithms
2:  Choose $(\alpha,\gamma_{1},\gamma_{2})$ such that $\alpha>0$, $0<\gamma_{2}\leq\gamma_{1}<\infty$.
@@
At $k$th iteration, the standard BB method employs the second-order model $m_{k}(p)=\nabla f(x_{k})^{T}p+\frac{1}{2}p^{T}H_{k}p$ where $H_{k}={\mu_{k}}^{-1}I$; thus the iterates have the form (2). Regarding the computation of the steplength, given $x_{k+1}$, the main idea of the BB method is to form the parameter $\mu_{k+1}^{\text{BB}}$ which solves
@@
In the BB method, $H_{k+1}=\mu_{k+1}^{-1}I$ is formed using $|\mu_{k+1}^{\text{BB}}|$ and thresholding $|\mu_{k+1}^{\text{BB}}|$ by means of two prefixed scalars $0<\mu_{min}<\mu_{max}$, i.e., setting $\mu_{k+1}=\max\{\mu_{min},\min\{|\mu_{k+1}^{\text{BB}}|,\mu_{max}\}$ to prevent excessively small or large values; in the case of convex problems it trivially holds $\mu_{k+1}=\max\{\mu_{min},\min\{\mu_{k+1}^{\text{BB}},\mu_{max}\}$. We refer to [27] for a survey on both practical rules for fixing $\mu_{k+1}$ and step acceptance rules in the deterministic convex and nonconvex settings.
@@
At the generic $k$-th iteration of TRishBB, the stochastic gradient $g_{k}$ is computed. Then, given the trust-region radius $\Delta_{k}$ and the scalar $\mu_{k}>0$, we set $H_{k}={\mu_{k}}^{-1}I$ and let (3) be the trust-region problem with strictly convex quadratic model. The trust-region solution is cheap to compute and takes the form
@@
When $p_{k}$ takes the form $p_{k}=-H_{k}^{-1}g_{k}=-\mu_{k}g_{k}$, it holds $\|p_{k}\|<\Delta_{k}$ and $p_{k}$ is the unconstrained minimizer of the model $m_{k}$. We refer to such a step as the unconstrained solution of (3). Otherwise, the solution lies on the boundary of the trust-region and it is the constrained Cauchy point; we will refer to this step as the constrained solution. In both cases, $p_{k}$ is a step along $(-g_{k})$ and the cost for solving the trust-region problem is low. On the contrary, using a second-order trust-region methods and a general $H_{k}$ requires a linear algebra phase such as the application of the conjugate gradient method [28, §7.5.1].
The updating rule (4) for $\Delta_{k}$ is a distinctive feature of TRish. It is ruled by three positive and prefixed scalars, $\alpha$, $\gamma_{1}$ and $\gamma_{2}$ and the magnitude of the norm of $g_{k}$. The insight of this rule is given in [7, §2]; briefly, adopting the standard rule, i.e., setting $\Delta_{k}=\alpha$, and thus normalizing $(-g_{k})$ if the trust-region solution is constrained, may cause the algorithm to compute ascent directions more likely than descent directions and then fail to progress in expectation. We note that in [7, 9] the algorithm admits the use of positive sequences $\{\alpha_{k}\}$, $\{\gamma_{1,k}\}$ and $\{\gamma_{2,k}\}$ but here we focus on the choice of constant parameters.
@@
$p_{k}=\begin{cases}-\alpha\gamma_{1}g_{k}&\quad\text{if }\|g_{k}\|\in\left[0,\frac{1}{\gamma_{1}}\right),\\ -\displaystyle\alpha\frac{g_{k}}{\|g_{k}\|}&\quad\text{if }\|g_{k}\|\in\left[\frac{1}{\gamma_{1}},\frac{1}{\gamma_{2}}\right],\\ -\alpha\gamma_{2}g_{k}&\quad\text{if }\|g_{k}\|\in\left(\frac{1}{\gamma_{2}},\infty\right),\end{cases}$ (8)
@@
$\mu_{k}<\begin{cases}\alpha\gamma_{1}&\quad\text{if }\|g_{k}\|\in\left[0,\frac{1}{\gamma_{1}}\right),\\ \displaystyle\frac{\alpha}{\|g_{k}\|}&\quad\text{if }\|g_{k}\|\in\left[\frac{1}{\gamma_{1}},\frac{1}{\gamma_{2}}\right],\\ \alpha\gamma_{2}&\quad\text{if }\|g_{k}\|\in\left(\frac{1}{\gamma_{2}},\infty\right).\end{cases}$ (9)
Importantly, once the step $p_{k}$ is formed, the new iterate $x_{k+1}$ is computed without performing the standard test of trust-region methods on predicted versus actual reduction [28]; thus the Algorithm TRishBB does not require the evaluation of the objective function $f$.
@@ 2.1 TRishBB_v1
The displayed Algorithm TRishBB_v1 is a straightforward implementation of the BB method employing stochastic gradients and guidelines from stochastic quasi-Newton methods. Specifically, once the trust-region solution $p_{k}$ in (7) is computed, the successive steplength $\mu_{k+1}$ is formed using the correction pair $(s_{k},y_{k})$ at line 5 of this algorithm. The vector $s_{k}$ coincides with $p_{k}$, i.e., it is the step between $x_{k+1}$ and $x_{k}$ whereas the vector $y_{k}$ is the difference between the mini-batch gradients $\nabla f_{\mathcal{N}_{k}}(x_{k+1})$ and $\nabla f_{\mathcal{N}_{k}}(x_{k})$. We underline that using the same set of indices $\mathcal{N}_{k}$ in the computation of $y_{k}$ is suggested in [29, 16, 17] and implies that (6) holds with the matrix $\int_{0}^{1}\nabla^{2}f_{\mathcal{N}_{k}}(x_{k}+ts_{k})$. Thus, the BB steplength inherits spectral information as long as it is positive and $f$ is twice continuously differentiable. Following a heuristic common rule, see e.g., [30], given an integer $m\geq 1$, the parameter $\mu_{k+1}$ is updated every $m$ iterations.
@@
1:  Given $x_{k}\in\mathbb{R}^{n}$, $(\alpha,\gamma_{1},\gamma_{2})$, $\mu_{k}>0$, $0<\mu_{min}<\mu_{max}$, $m\geq 1$, $\beta=\frac{m-1}{m}$, $\eta\in(0,1)$, $\bar{x}_{old}$, $\bar{g}_{old}$, and $\bar{g}$  (if $k=0$, then $\bar{x}_{old}=x_{0}$, $\bar{g}_{old}=0$, $\bar{g}=0$, and $\bar{\mu}_{0}={\mu}_{0}$):
@@
6:     $\bar{x}=x_{k+1}$;
7:     $s_{k}=\bar{x}-\bar{x}_{old}$ and $y_{k}=\bar{g}-\bar{g}_{old}$;
8:      $\hat{\mu}_{k}=\frac{1}{m}\left|\frac{s_{k}^{T}s_{k}}{s_{k}^{T}y_{k}}\right|$, $\,\bar{\mu}_{k+1}=\eta\bar{\mu}_{k}+(1-\eta)\hat{\mu}_{k}$;
9:      $\mu_{k+1}=\max\{\mu_{min},\min\{\bar{\mu}_{k+1},\mu_{max}\}\}$;
10:     $\bar{x}_{old}=\bar{x}$, $\bar{g}_{old}=\bar{g}$, and $\bar{g}=0$;
@@ 2.2 TRishBB_v2
The vector $s_{k}$ is the difference between two, not consecutive, iterates; the indices of such iterates differ by $m$, see lines 7 and 10 of the algorithm. The vector $y_{k}$ is the difference between vectors computed via a moving average rule which accumulates stochastic gradients over $m$ iterations, see lines 4, 7 and 10 of the algorithm; in particular $\bar{g}$ is formed over the last $m$ iterations while $\bar{g}_{old}$ is the analogous vector formed in the previous cycle of $m$ iterations. The parameter $\beta=(m-1)/m$ is fixed in accordance with [31] to avoid hyper-parameter tuning.
The form of $s_{k}$ and $y_{k}$ is borrowed from the BB-SGD algorithm [11] where $|\mathcal{N}_{k}|=1$, $\forall k$, and $m=N$, i.e., $N$ single gradient evaluations are performed before updating the steplength. In TRishBB_v2 we allow $|\mathcal{N}_{k}|>1$; letting $N_{b}=\lfloor N/|\mathcal{N}_{k}|\rfloor$ be (roughly) the number of disjoint mini-batches of size $|\mathcal{N}_{k}|$ in $\mathcal{N}$ and $m=N_{b}$, the cost for computing $\bar{g}$ every $m$ iterations is (roughly) the cost for one full gradient evaluation.
The steplength $\mu_{k+1}$ is updated every $m$ iterations and computed using accumulated information, see lines 8 and 9 of the algorithm. Given $\eta\in(0,1)$, the scalar $\bar{\mu}_{k+1}$ is computed as the convex combination of the current stochastic BB steplength $\hat{\mu}_{k}$ and the current accumulated parameter $\bar{\mu}_{k}$. The larger $\eta$ is, the higher the penalization to $\hat{\mu}_{k}$ is and this may soften the effect of intrinsic noise arising from stochastic gradients.
@@
1:  Given $x_{k}\in\mathbb{R}^{n}$, $(\alpha,\gamma_{1},\gamma_{2})$, $\mu_{k}>0$ with $0<\mu_{min}<\mu_{max}$, $m\geq 1$, $L\in\mathbb{N}$ $F_{k}\in\mathbb{R}^{n\times d}$ with $d\leq L$, $x_{avg}$, and $\bar{x}_{old}$,   (if $k=0$, then $F_{0}=[\,]$, $x_{avg}=0$, and $\bar{x}_{old}=0$).
2:  $x_{avg}=x_{avg}+x_{k}$
@@
5:  if $k\leq L$ then
6:     Store $g_{k}$ in $k$th column of $F_{k}$.
@@
8:     Remove the first column of $F_{k}$, and store $g_{k}$ in its $L$-th column.
@@
14:        $\mu_{k+1}=\max\left\{\mu_{min},\min\left\{\left|\frac{{s}_{k}^{T}{s}_{k}}{{s}_{k}^{T}{y}_{k}}\right|,\mu_{max}\right\}\right\}$
15:     end if
16:     $\bar{x}_{old}=\bar{x}$
17:  else
18:     $\mu_{k+1}=\mu_{k}$
19:  end if

### v2

@@ Abstract
We investigate stochastic gradient methods and stochastic counterparts of the Barzilai-Borwein steplengths and their application to finite-sum minimization problems. Our proposal is based on the Trust-Region-ish (TRish) framework introduced in [F. E. Curtis, K. Scheinberg, R. Shi, A stochastic trust region algorithm based on careful step normalization, Informs Journal on Optimization, 1, 2019]. The new framework, named TRishBB, aims to enhance the performance of TRish and at reducing the computational cost of the second-order TRish variant. We propose three different methods belonging to the TRishBB framework and present the convergence analysis for possibly nonconvex objective functions, considering biased and unbiased gradient approximations. Our analysis requires neither diminishing step-sizes nor full gradient evaluation. The numerical experiments in machine learning applications demonstrate the effectiveness of applying the Barzilai-Borwein steplength with stochastic gradients and show improved testing accuracy compared to the TRish method.
@@ 1 Introduction
Computing the gradient $\nabla f(x_{k})$ to perform the update (2) is practically inconvenient in ML and DL applications where $f$ represents a loss function on a given dataset and $N$ is large. The Stochastic Gradient (SG) method [4, 5] and the mini-batch SG method [6] are iterative procedures where the full gradient $\nabla f(x_{k})$ is replaced by either a single gradient $\nabla f_{i}(x_{k})$, $i\in\mathcal{N}$ selected at random, or by a mini-batch of samples of the form $\nabla f_{\mathcal{N}_{k}}(x_{k})=\frac{1}{|\mathcal{N}_{k}|}\sum_{i\in\mathcal{N}_{k}}\nabla f_{i}(x_{k})$ where the set $\mathcal{N}_{k}\subseteq\mathcal{N}$ is selected at random and $|\mathcal{N}_{k}|$ denotes its cardinality, i.e., the mini-batch size. A limitation of such methods is the need to choose problem-dependent steplengths to guarantee convergence. An attempt to reduce such dependence was made in [7] where the steplengths depend on a Trust-Region-ish (TRish) rule and normalized steps are selected whenever the norm of the stochastic gradient is within a prefixed interval whose endpoints need to be tuned. The TRish method was designed for both first-order models [7, 8] and quadratic models [9] and, contrary to a standard trust-region approach, does not test for actual vs. predicted reduction and thus does not call for $f$-evaluations. Results from [7, 8] show that TRish can outperform a standard stochastic gradient approach and it was recently extended to equality-constrained problems [10].
@@
This paper is organized as follows. The remainder of this section is devoted to a review of the relevant literature providing background, motivation, and contribution to our work. We outline the proposed stochastic algorithms in Section 2 and present the theoretical analysis in Section 3. In Section 4 we present the implementation details, experimental setup, and an extensive evaluation of the method’s performance. Finally, in Section 5 we provide an overview of the obtained results and perspectives.
@@ 1.1 Related works and motivations
In [11], BB steplengths were incorporated into the Stochastic Variance Reduction Gradient (SVRG) method [23] and into the SG method. The proposed methods are inner-outer loop-based frameworks. Since SVRG provides a full gradient at each outer loop, the BB steplength was computed at outer loops using the standard formula for the BB steplength; convergence of the resulting procedure was proved for strongly convex objective functions. On the other hand, single stochastic gradients $\nabla f_{i}$ were employed at each iteration in the SG method and, in the inner loop, a moving averaging rule over $N$ stochastic gradients at different points was computed; then, such average was used at each outer loop to evaluate a stochastic BB parameter. A theoretical support for this algorithm, denoted SGD-BB algorithm, was not provided. In [12], deterministic BB steplengths from [24], based on exact gradient and function evaluations, were used in the SVRG method; convergence was analyzed for strongly convex objective functions. In [13], stochastic BB steplengths were formed by a moving average rule on mini-batch gradients evaluated at different points and selected if they lied within an interval with endpoints tending to zero. Using diminishing stepsizes provides convergence due to well-known results for SG method but, in principle, BB steplengths may be discarded eventually since they are not supposed to tend to zero. A further method in which stochastic BB steplengths were formed and diminishing stepsizes were imposed is given in [14]; the distinguishing feature is that the mini-batch of sample is taken constant for a prefixed number of iterations with the aim to explore the underlying approximate Hessian spectrum. Finally, the paper [15] investigates on the use of BB steplengths combined with projected stochastic gradients for constrained optimization problems and with the objective function in the form of mathematical expectation.
@@ 2 New Stochastic Trust-Region-ish Algorithms
2: Choose $(\alpha,\gamma_{1},\gamma_{2})$ such that $\alpha>0$, $0<\gamma_{2}\leq\gamma_{1}<\infty$ and $\mu_{min}>0$.
@@
At $k$-th iteration, the standard BB method minimizes the second-order model $m_{k}(p)=\nabla f(x_{k})^{T}p+\frac{1}{2}p^{T}H_{k}p$ where $H_{k}=\mu_{k}^{-1}I$ for some positive $\mu_{k}$; thus the iterates have the form (2). Regarding the computation of the steplength, given $x_{k+1}$, the main idea of the BB method is to form the parameter $\mu_{k+1}^{\text{BB}}$ which solves
@@
In the BB method, $H_{k+1}=\mu_{k+1}^{-1}I$ is formed using $|\mu_{k+1}^{\text{BB}}|$ and thresholding $|\mu_{k+1}^{\text{BB}}|$ by means of two predefined scalars $0<\mu_{min}<\mu_{max}$, i.e., setting $\mu_{k+1}=\max\{\mu_{min},\min\{|\mu_{k+1}^{\text{BB}}|,\mu_{max}\}$ to prevent excessively small or large values; in the case of convex problems it trivially holds $\mu_{k+1}=\max\{\mu_{min},\min\{\mu_{k+1}^{\text{BB}},\mu_{max}\}$. We refer to [27] for a survey on both practical rules for fixing $\mu_{k+1}$ and step acceptance rules in the deterministic convex and nonconvex settings.
@@
At the generic $k$-th iteration of TRishBB, the stochastic gradient $g_{k}$ is computed. Then, given the trust-region radius $\Delta_{k}$ and the scalar $\mu_{k}>\mu_{min}$, we set $H_{k}=\mu_{k}^{-1}I$ and let (3) be the trust-region problem with strictly convex quadratic model. The trust-region solution is cheap to compute and takes the form
@@
In fact, when $p_{k}$ takes the form $p_{k}=-H_{k}^{-1}g_{k}=-\mu_{k}g_{k}$, it holds $\|p_{k}\|<\Delta_{k}$ and $p_{k}$ is the unconstrained minimizer of the model $m_{k}$. We refer to such a step as the unconstrained solution of (3). Otherwise, the solution lies on the boundary of the trust-region and it is the constrained Cauchy point; we will refer to this step as the constrained solution. In both cases, $p_{k}$ is a step along $(-g_{k})$ and the cost for solving the trust-region problem is low. On the contrary, using a second-order trust-region methods and a general $H_{k}$ requires a linear algebra phase such as the application of the conjugate gradient method [28, §7.5.1].
The updating rule (4) for $\Delta_{k}$ is a distinctive feature of TRish. It is ruled by three positive and predefined scalars, $\alpha$, $\gamma_{1}$ and $\gamma_{2}$ and the magnitude of the norm of $g_{k}$. The insight of this rule is given in [7, §2]; briefly, adopting the standard rule, i.e., setting $\Delta_{k}=\alpha$, and thus normalizing $(-g_{k})$ if the trust-region solution is constrained, may cause the algorithm to compute ascent directions more likely than descent directions and then fail to progress in expectation. We note that in [7, 9] the algorithm admits the use of positive sequences $\{\alpha_{k}\}$, $\{\gamma_{1,k}\}$ and $\{\gamma_{2,k}\}$ but here we focus on the choice of constant parameters.
@@
$p_{k}=\begin{cases}-\alpha\gamma_{1}g_{k}&\quad\text{if }\|g_{k}\|\in\left[0,\frac{1}{\gamma_{1}}\right)\\ -\displaystyle\alpha\frac{g_{k}}{\|g_{k}\|}&\quad\text{if }\|g_{k}\|\in\left[\frac{1}{\gamma_{1}},\frac{1}{\gamma_{2}}\right]\\ -\alpha\gamma_{2}g_{k}&\quad\text{if }\|g_{k}\|\in\left(\frac{1}{\gamma_{2}},\infty\right)\end{cases},$ (8)
@@
$\mu_{k}<\begin{cases}\alpha\gamma_{1}&\quad\text{if }\|g_{k}\|\in\left[0,\frac{1}{\gamma_{1}}\right)\\ \displaystyle\frac{\alpha}{\|g_{k}\|}&\quad\text{if }\|g_{k}\|\in\left[\frac{1}{\gamma_{1}},\frac{1}{\gamma_{2}}\right]\\ \alpha\gamma_{2}&\quad\text{if }\|g_{k}\|\in\left(\frac{1}{\gamma_{2}},\infty\right)\end{cases}.$ (9)
Importantly, once the step $p_{k}$ is formed, the new iterate $x_{k+1}$ is computed without performing the standard test of trust-region methods on predicted vs. actual reduction [28]; thus the Algorithm TRishBB does not require the evaluation of the objective function $f$.
@@ 2.1 TRishBB_v1
The displayed Algorithm TRishBB_v1 is a straightforward implementation of the BB method employing stochastic gradients and guidelines from stochastic quasi-Newton methods. Specifically, once the trust-region solution $p_{k}$ in (7) is computed, the successive steplength $\mu_{k+1}$ is formed using the correction pair $(s_{k},y_{k})$ at line 5 of this algorithm. The vector $s_{k}$ coincides with $p_{k}$, i.e., it is the step between $x_{k+1}$ and $x_{k}$ whereas the vector $y_{k}$ is the difference between the mini-batch gradients $\nabla f_{\mathcal{N}_{k}}(x_{k+1})$ and $\nabla f_{\mathcal{N}_{k}}(x_{k})$. We underline that using the same set of indexes $\mathcal{N}_{k}$ in the computation of $y_{k}$ is suggested in [29, 16, 17] and implies that (6) holds with the matrix $\int_{0}^{1}\nabla^{2}f_{\mathcal{N}_{k}}(x_{k}+ts_{k})$. Thus, the BB steplength inherits spectral information as long as it is positive and $f$ is twice continuously differentiable. Following a heuristic common rule, see e.g., [30], given an integer $m\geq 1$, the parameter $\mu_{k+1}$ is updated every $m$ iterations.
Compared to TRish, the extra cost of this algorithm amounts to the computation of $\nabla f_{\mathcal{N}_{k}}(x_{k+1})$ and two scalar products to evaluate $\mu_{k+1}$.
@@
1: Given $x_{k}\in\mathbb{R}^{n}$, $(\alpha,\gamma_{1},\gamma_{2})$, $\mu_{k}>0$, $\bar{\mu}>0$, $0<\mu_{min}<\mu_{max}$, $m\geq 1$, $\beta=\frac{m-1}{m}$, $\eta\in(0,1)$, $\bar{x}_{old}$, $\bar{g}_{old}$, $\bar{g}$  (if $k=0$, then $\bar{x}_{old}=x_{0}$, $\bar{g}_{old}=0$, $\bar{g}=0$, and $\bar{\mu}={\mu}_{0}$):
@@
6:  $\bar{x}=x_{k+1}$
7:  $s_{k}=\bar{x}-\bar{x}_{old}$ and $y_{k}=\bar{g}-\bar{g}_{old}$
8:   $\widehat{\mu}_{k}=\frac{1}{m}\left|\frac{s_{k}^{T}s_{k}}{s_{k}^{T}y_{k}}\right|$, and $\bar{\mu}=\eta\bar{\mu}+(1-\eta)\widehat{\mu}_{k}$
9:   $\mu_{k+1}=\max\{\mu_{min},\min\{\bar{\mu},\mu_{max}\}\}$
10:  $\bar{x}_{old}=\bar{x}$, $\bar{g}_{old}=\bar{g}$, $\bar{g}=0$
@@ 2.2 TRishBB_v2
The vector $s_{k}$ is the difference between two, not consecutive, iterates; the indexes of such iterates differ by $m$, see lines 7 and 10 of the algorithm. The vector $y_{k}$ is the difference between vectors computed via a moving average rule which accumulates stochastic gradients over $m$ iterations, see lines 4, 7 and 10 of the algorithm; in particular $\bar{g}$ is formed over the last $m$ iterations while $\bar{g}_{old}$ is the analogous vector formed in the previous cycle of $m$ iterations. The parameter $\beta=(m-1)/m$ is fixed in accordance with [31] to avoid hyper-parameter tuning.
The form of $s_{k}$ and $y_{k}$ and $\widehat{\mu}_{k}$ is borrowed from the stochastic gradient descent algorithm with BB step size (SGD-BB) algorithm proposed in [11] where $|\mathcal{N}_{k}|=1$, $\forall k$, and $m=N$, i.e., $N$ single gradient evaluations are performed before updating the steplength. In TRishBB_v2 we allow $|\mathcal{N}_{k}|>1$; letting $N_{b}=\lfloor N/|\mathcal{N}_{k}|\rfloor$ be (roughly) the number of disjoint mini-batches of size $|\mathcal{N}_{k}|$ in $\mathcal{N}$ and $m=N_{b}$, the cost for computing $\bar{g}$ through $m$ iterations is approximately the cost for one full gradient evaluation.
The steplength $\mu_{k+1}$ is updated every $m$ iterations and computed using accumulated information, see lines 8 and 9 of the algorithm. Given $\eta\in(0,1)$, the scalar $\bar{\mu}$ is computed as the convex combination of the current stochastic BB steplength $\widehat{\mu}_{k}$ and the current accumulated parameter $\bar{\mu}$. The larger $\eta$ is, the higher the penalization to $\widehat{\mu}_{k}$ is and this may soften the effect of intrinsic noise arising from stochastic gradients.
Compared to TRish, the extra cost of this algorithm amounts to two scalar products to evaluate $\mu_{k+1}$; in terms of storage, three vectors of size $n$ are required.
@@
1: Given $x_{k}\in\mathbb{R}^{n}$, $(\alpha,\gamma_{1},\gamma_{2})$, $\mu_{k}>0$, $\bar{\mu}>0$, $0<\mu_{min}<\mu_{max}$, $m\geq 1$, $m_{F}\in\mathbb{N}$ $F_{k}\in\mathbb{R}^{n\times q}$ with $q\leq m_{F}$, $x_{avg}$, and $\bar{x}_{old}$,   (if $k=0$, then $F_{0}=[\,]$, $x_{avg}=0$, and $\bar{x}_{old}=0$, and $\bar{\mu}={\mu}_{0}$).
[…]
@@
5:if $k\leq m_{F}$ then
6:  Store $g_{k}$ in $k$-th column of $F_{k}$.
@@
8:  Remove the first column of $F_{k}$, and store $g_{k}$ in its $m_{F}$-th column.
@@
14:    $\widehat{\mu}_{k}=\frac{1}{m}\left|\frac{s_{k}^{T}s_{k}}{s_{k}^{T}y_{k}}\right|$, and $\bar{\mu}=\eta\bar{\mu}+(1-\eta)\widehat{\mu}_{k}$;
15:    $\mu_{k+1}=\max\{\mu_{min},\min\{\bar{\mu},\mu_{max}\}\}$;
16:  end if
17:  $\bar{x}_{old}=\bar{x}$
18:else
19:  $\mu_{k+1}=\mu_{k}$
20:end if